the volume of the box is 560 inches. the width is 7 inches. how long and heigh of the box

Answer by checkley77(12844)

W*L*H=560
7*LH=560
LH=560/7
LH=80
THUS THE PRODUCT OF THE LENGTH*HEIGHT MUST=80.
ONE COMBINATION WOULD BE LENGTH=8 & HEIGHT=10 8*10=80.
OR:
LENGTH=20 & HEIGHT=4 20*4=80.